HOT-132: Chief Scientist Report
Chief Scientist: J. JOHNSON
HOT-132 Chief Scientist's Cruise Report
R/V Ka'Imikai O Kanaloa
November 15-19, 2001
Departed: November 15, 2001 at 0900 (HST)
Returned: November 19, 2001 at 0700
Vessel: R/V Ka'Imikai O Kanaloa
Operator: University of Hawaii
Master of the Vessel: Captain Robert Hayes
Chief Scientist: Jeremiah Johnson
STAG Electronics Technician: Steve Poulos
STAG Deck Operations: David Gravatt
1. SCIENTIFIC OBJECTIVES
The objective of this cruise was to maintain a collection of
hydrographic and biogeochemical data at the Hawaii Ocean Time-series
(HOT) stations. Four stations were to be occupied during the cruise,
in the following order:
1) Station 1, referred to as Station Kahe, is located at 21o 20.6'N,
158o 16.4'W and was to be occupied on November 15 for about 3 hours.
2) Station 2: ALOHA (A Long Term Oligotrophic Habitat Assessment) is
defined as a circle with a 6 nautical mile radius centered at 22o 45'N,
158oW. This is the main HOT station and was to be occupied for 2 days
from November 16 and 17.
3) Station 8, referred to as HALE-ALOHA is the location of the deep
ocean mooring (22o 20.0'N, 158o 10.6'W). It was to be occupied on
November 18 for about 2 hours.
4) Station 6, referred to as Station Kaena , is located off Kaena Point
at 21 50.8'N, 158 21.8'W was to be occupied on November 18 for about 3 hours.
A single CTD cast was to be conducted at Station 1 to collect
continuous profiles of various physical and chemical parameters. Water
samples were to be collected at discrete depths for biogeochemical
Upon arrival at Station ALOHA, a net tow was to be conducted, followed
by the deployment of a free-drifting sediment trap array. After
deployment, a full-depth CTD cast was to be conducted, followed by CTD
casts at strict 3 hour intervals for at least 36 hours for continuous and
discrete data collection, followed by another full-depth CTD cast. Two
other free-drifting arrays were to be deployed on November 17: an oxygen
balance experiment (O2) for 24 hours, and a primary production experiment
for 12 hours. A plankton net was to be deployed near noon and midnight on
November 16 and 17 at Station ALOHA.
After work at Station ALOHA was accomplished, the ship was to transit to
recover the sediment trap array. After the sediment traps were recovered,
the ship was to transit to Station 8, to conduct one
1000-m CTD cast, after which the ship was to transit to station 6.
A near-bottom CTD cast (~2500 m) was to be conducted at Station 6
including salinity samples for calibration, after which the ship was to
transit back to Snug Harbor.
A Profiling Reflectance Radiometer (PRR) and a Tethered Spectral
Radiometer Buoy (TSRB) was to be deployed for half-hour periods near
noon time on November 15, 16 and 17.
The following instruments were to collect data throughout the cruise: a
shipboard ADCP, a thermosalinograph and fluorometer, and an anemometer.
2. SCIENCE PERSONNEL
Jeremiah Johnson (Cheif Scientist) Research Associate UH
Noel Larson (Watch leader) Research Associate UH
John Hargrove Volunteer UH
Tara Clemente Research Associate UH
Fernando Santiago-Mandujano Research Associate UH
Mark Valenciano Electronics Technician UH
Cecilia Sheridan Graduate Student UH
Mya Iriondo-Simek Graduate Student UH
Karin Bjorkman Scientist UH
Lance Fujieki Computer Specialist UH
Anne Gasc Research Associate UH
Dale Hebel Scientist UH
Paul Morris Technician UH
Dan Drown graduate Student UH
Tom Gregory (Watch Leader) Research Associate UH
Rodrigo Gonzales Scientist WHOI
3. GENERAL SUMMARY
All scientific objectives were met. Thirteen 1000-m CTD casts and two
deep casts were obtained at station ALOHA. One 1000-m CTD cast was
obtained at station Kahe and one at HALE-ALOHA. One near-bottom
cast(~2500 m) was obtained at station 6.
The array of floating sediment traps, the primary productivity array,
and P. Morris' O2 array were deployed and recovered without incidents.The
sediment traps array drifted 20 nm North from Station ALOHA.
C. Sheridan completed successfully 6 pairs of plankton net tows.
The PRR and TSRB we successfully deployed.
Winds were light and variable.
The ADCP ran without interruption throughout the cruise, as well as
thethermosalinograph, fluorometer, and the ship's anemometer.
We arrived back at Snug Harbor on November 19 at 0700. Complete off-load
took place immediately.
4. R/V KA'IMIKAI O KANALOA, OFFICERS AND CREW, TECHNICAL SUPPORT
The R/V Ka'Imikai O Kanaloa continues to maintain the excellent ship
support for our work. The officers and crew were most helpful and
accommodating. They showed enthusiasm and concern for our work and
were very flexible in receiving changes in our operational schedule.
Technical support during this cruise was excellent. STAG personnel
wereavailable at any time to assist in our work and made things much
5. DAILY REPORT OF ACTIVITIES (HST)
November 14, 2001; Loading Day
Equipment loaded on this day. Tested CTD system.
November 15, 2001
The ship departed from Snug harbor at 0900. Fire and abandon ship drills
conducted at 0930, followed by a short science meeting during which the
cruise schedule was reviewed, and safety issues were addressed.
Arrived to Kahe station at 1130, and a weight cast (400 lb) to 1000m
was conducted during which M. Valenciano inspected the CTD wire. At
1300 the Profiling Reflectance Radiometer (PRR) and the Tethered Spectral
Radiometer Buoy (TSRB)were deployed.
A CTD cast was conducted at 1400. We then transited to ALOHA.
Winds were light from the east.
November 16, 2001
Arrived at station ALOHA at 23:30 (Nov 15). We then did a net tow and
then deployed the sediment traps. The deep WOCE cast started at 0200.
The STAG fluorometer failed on the deep cast. We used the JGOFS
fluorometer for the remainder of the cruise. The primary oxygen sensor
failed during the WOCE deep cast. It was replaced and the replacement
failed on s2c2. We then installed another oxygen sensor on the primary
channel which functioned properly for the rest of the cruise. The
seconday oxygen sensor worked for the entire cruise. The 36-hr CTD cast
period started at 0800. A total of six 1000-m CTD casts were conducted
Two pairs of net tows were conducted during the day and another pair at
At 1200 the Profiling Reflectance Radiometer (PRR) and the Tethered
Spectral Radiometer Buoy (TSRB) were deployed.
Winds were light from the east with a large ground-swell.
November 17, 2001
Seven 1000-m CTD casts were conducted during this day, ending the 36-hr
CTD cast period at 2100. The WOCE deep cast was started at 2300.
The O2 array was deployed at 0400. The primary productivity array was
deployed at 0600 and was retrieved at 1800. The PP array drifted 10nm
north of the circles center.
One pair of net tows was conducted in the day and night.
At 1200 the Profiling Reflectance Radiometer (PRR) and the Tethered
Spectral Radiometer Buoy (TSRB)were deployed.
Winds were light from the southeast, with a large ground-swell.
November 18, 2001
The deep CTD cast that started at 2300 on November 17 was completed by
0300 on November 18.
The O2 array was recovered at 0630.
The sediment trap array was recovered at 0900. The array drifted 20 nm
North from Station ALOHA.
A 1000-m CTD cast was conducted at the HALE-ALOHA station at 1400.
A near-bottom CTD cast (~2500 m) was conducted at 1900 at station 6.
Winds were light from the south with a little less groundswell.
November 19, 2001
Arrived at Snug Harbor at 0700. Complete off-load took place immediately.
Sub component programs:
Bob Bidigare HPLC pigments/UH
Michael Landry zooplankton dynamics/UH
John Dore CO2 dynamics/UH
Charles Keeling CO2 dynamics and intercalibration/SIO
Paul Quay DIC and 13C/UW
Abbott/Letelier optical measurements/OSU
Claudia Benitez-Nelson phosphorus isotopes, Th234/UH
Peter J. LeB. Williams oxygen balance/U Wales Bangor
Hebel, Dore, Karl EOC/UH
John Dore P15N/UH
Paul Morris oxygen balance/UH
Karin Bjorkman Phosphorus dynamics/UH
R. Gonzales Electron transfer system
P. Morris, T. Gregory